# references.bib

### v1

@@
Abstract
We compute the Hilbert series of the space of $n=3$ variable quasi-invariant polynomials in characteristic $2$ and $3,$ capturing the dimension of the homogeneous components of the space, and explicitly describe the generators in the characteristic $2$ case. In doing so we extend the work of the first author in 2023 on quasi-invariant polynomials in characteristic $p>n$ and prove that a sufficient condition found by Ren–Xu in 2020 on when the Hilbert series differs between characteristic 00 and $p$ is also necessary for $n=3,p=2,3$. This is the first description of quasi-invariant polynomials in the case where the space forms a modular representation over the symmetric group, bringing us closer to describing the quasi-invariant polynomials in all characteristics and numbers of variables.
@@ 1 Introduction
Let $\mathbbm{k}$ be a field, and consider the action of the symmetric group $S_{n}$ on the space $\mathbbm{k}[x_{1},\dots,x_{n}]$ of $\mathbbm{k}$-valued polynomials by permuting the variables. A polynomial in $\mathbbm{k}[x_{1},\dots,x_{n}]$ is symmetric if it is invariant under this action. Equivalently, since $S_{n}$ is generated by transpositions, a polynomial $K$ is symmetric if $s_{ij}K=K$ or $(1-s_{ij})K=0$ for all $s_{ij}\in S_{n}$. One may consider generalizations of symmetric polynomials in which this condition is relaxed, so that we only require $(1-s_{ij})K$ be divisible by some large polynomial. This leads to the notion of quasi-invariant polynomials:
@@
Let $\mathbbm{k}$ be a field. For $m\in\mathbb{Z}_{\geq 0},n\in\mathbb{Z}_{>0}$, a polynomial $K\in\mathbbm{k}[x_{1},\dots,x_{n}]$ is $m$-quasi-invariant if for all $s_{ij}\in S_{n}$ we have that $(x_{i}-x_{j})^{2m+1}$ divides $(1-s_{ij})K$. We denote the space of $m$-quasi-invariants by $Q_{m}(n,\mathbbm{k})$.
@@
Quasi-invariant polynomials were first introduced by Chalykh and Veselov in 1990 [cmp/1104179957] to describe the harmonic, zero eigenvalue eigenfunctions of quantum Calogero-Moser systems. Calogero-Moser systems are a collection of one-dimensional dynamical particle systems that were found to be both solvable [calogero] and integrable [MOSER1975197]. Due to these properties, they have become extensively studied in mathematical physics, with connections to a number of other fields of mathematics, including representation theory.
Quasi-invariant polynomials were also later found to describe the representation theory of the spherical subalgebra of the rational Cherednik algebra [beg]. This subalgebra is Morita equivalent to the entire rational Cherednik algebra [Etingof2000SymplecticRA], so quasi-invariants describe representations of rational Cherednik algebras as well. Such algebras have connections to combinatorics, mathematical physics, algebraic geometry, algebraic topology, and more, leading them to become a central topic in representation theory.
@@
In 2003, Felder and Veselov found the Hilbert series of the space of quasi-invariants in characteristic zero [felder2001action], proving its freeness in the process. Work on quasi-invariants in characteristic $p$ started in 2020, when Ren and Xu proved a sufficient condition for the Hilbert series of $Q_{m}(n,\mathbf{F}_{p})$ to be different from the Hilbert series of $Q_{m}(n,\mathbf{Q})$ [Ren_2020]. They accomplished this by computing non-symmetric polynomial “counterexamples” in characteristic $p$ where the polynomial has lower degree than any non-symmetric quasi-invariant polynomial in characteristic 0. They also made several conjectures about quasi-invariants in characteristic $p$, including that the condition they found is also sufficient, the quasi-invariants are free, and that the Hilbert polynomial is palindromic for $p>2$. In 2023, the first author proved a general form for the Hilbert series of the quasi-invariants for $n=3,p>3$, proving freeness and the palindromicity of the Hilbert polynomial in the process [wang2022explicit].
In this paper, we consider the cases $n=3,p=2,3$. These cases differ from the $p>3$ case studied in [wang2022explicit] since in $p=2,3$ the representations of $S_{3}$ are modular, i.e. are not completely reducible. Despite these limitations, we describe the Hilbert series explicitly for all $m$, proving the following:
@@
$\mathcal{H}(Q_{m}(3,\mathbbm{k}))=\frac{1+2t^{d}+2t^{6m+3-d}+t^{6m+3}}{(1-t)(1-t^{2})(1-t^{3})}$
where $d=3m+1$ if there is no Ren–Xu counterexample and $d$ is the degree of the minimal degree Ren–Xu counterexample otherwise. In particular, the conditions found in [Ren_2020] for the Hilbert series of $Q_{m}(3,\mathbbm{k})$ to be different from the Hilbert series of $Q_{m}(3,\mathbf{Q})$ are necessary.
@@
In the case $p=2$, we also define $m$-quasi-invariants in the case where $m$ is a half-integer and prove an analogous statement to Theorem 1.3 in this case. Using quasi-invariants at half-integers, we also compute the generators of $Q_{m}(3,\mathbf{F}_{2})$ as a $\mathbf{F}_{2}[x_{1},x_{2},x_{3}]^{S_{3}}$-module explicitly.
In Section 2, we state some of the basic facts about quasi-invariant polynomials and introduce modular representations of $S_{3}$. In Section 3, we compute the generators of $Q_{m}(3,\mathbf{F}_{2})$, proving Theorem 1.3 for $p=2$ in the process. In Section 4, we begin discussing $p=3$, and show that some properties of quasi-invariants in 3 variables from [wang2022explicit] carry over to the $p=3$ case after converting from the standard representation to the $\mathrm{sign-triv}$ representation. In Section 5, we show that minimal degree Ren–Xu counterexamples are the lowest degree non-symmetric generators for $Q_{m}(3,\mathbf{F}_{3})$ and show that there is one other higher degree generator belonging to the $\mathrm{sign-triv}$ representation. Finally, in Section 6 we consider all other indecomposable representations of $S_{3}$ in $Q_{m}(3,\mathbf{F}_{3})$, finishing the proof of Theorem 1.3 for $p=3$.
@@ 2 Preliminaries
  1. 1)
    $\mathbbm{k}[x_{1},x_{2},x_{3}]^{S_{3}}\subset Q_{m}(3,\mathbbm{k})$, $Q_{0}(3,\mathbbm{k})=\mathbbm{k}[x_{1},x_{2},x_{3}],$ and $Q_{m}(3,\mathbbm{k})\supset Q_{m^{\prime}}(3,\mathbbm{k})$ where $m^{\prime}>m$.
  2. 2)
@@
We consider $Q_{m}(3,\mathbf{F}_{2})$ and $Q_{m}(3,\mathbf{F}_{3})$ as representations of $S_{3}$ where $S_{3}$ permutes the variables $x_{1},x_{2},x_{3}$. Since $Q_{m}(3,\mathbf{F}_{2})$ and $Q_{m}(3,\mathbf{F}_{3})$ are vector spaces over $\mathbf{F}_{2}$ and $\mathbf{F}_{3}$ respectively and the characteristics $2$ and $3$ divide $|S_{3}|$, $Q_{m}(3,\mathbf{F}_{2})$ and $Q_{m}(3,\mathbf{F}_{3})$ are modular representations of $S_{3}$.
@@ 2.1 Preliminary definitions for p=2
  2. 2)
    $\mathrm{std}$ is the $2$ dimensional irreducible representation of $S_{3}$ obtained by reducing the standard representation in characteristic 0 mod 2.
  3. 3)
    $\mathrm{triv-triv}$ is the $2$ dimensional indecomposable representation that contains a copy of $\mathrm{triv}$ as a subrepresentation such that the quotient of $\mathrm{triv-triv}$ by this subrepresentation is $\mathrm{triv}$.
@@
The polynomial $E_{\mathrm{triv-triv}}:=x_{1}^{2}x_{2}+x_{2}^{2}x_{3}+x_{3}^{2}x_{1}\in\mathbf{F}_{2}[x_{1},x_{2},x_{3}]$ generates a copy of $\mathrm{triv-triv}$. To see this, note that for any $i,j$, we have
$(1-s_{ij})E_{\mathrm{triv-triv}}=x_{1}^{2}x_{2}+x_{1}x_{2}^{2}+x_{1}^{2}x_{3}+x_{1}x_{3}^{2}+x_{2}^{2}x_{3}+x_{2}x_{3}^{2}\in\mathbf{F}_{2}[x_{1},x_{2},x_{3}]^{S_{3}}.$
@@
We cannot define $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv-triv}}$ to exclude copies of $\mathrm{triv}$ since we can add elements of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}}$ to copies of $\mathrm{triv-triv}$ and still obtain a copy of $\mathrm{triv-triv}.$ For example, $F:=E_{\mathrm{triv-triv}}+x_{1}^{3}+x_{2}^{3}+x_{3}^{3}$ still satisfies $(1-s_{ij})F=(1-s_{ij})E_{\mathrm{triv-triv}}$ for all $i,j$, so it generates a copy of $\mathrm{triv-triv}$ by the same argument as Example 2.4.

### v2

# Hilbert Series of $S_{3}$-Quasi-Invariant Polynomials in Characteristics 2, 3

@@
[…]
Hilbert Series of $S_{3}$-Quasi-Invariant Polynomials in Characteristics 2, 3
[…]
Hilbert Series of $\boldsymbol{S_{3}}$-Quasi-Invariant Polynomials
in Characteristics 2, 3
[…]
Frank WANG and Eric YEE
[…]
F. Wang and E. Yee
[…]
Department of Mathematics, Massachusetts Institute of Technology,
[…]
Received January 31, 2025, in final form July 02, 2025; Published online July 13, 2025
[…]
We compute the Hilbert series of the space of $n=3$ variable quasi-invariant polynomials in characteristic $2$ and $3$, capturing the dimension of the homogeneous components of the space, and explicitly describe the generators in the characteristic 2 case. In doing so we extend the work of the first author in 2023 on quasi-invariant polynomials in characteristic $p>n$ and prove that a sufficient condition found by Ren–Xu in 2020 on when the Hilbert series differs between characteristic 00 and $p$ is also necessary for $n=3$, $p=2,3$. This is the first description of quasi-invariant polynomials in the case, where the space forms a modular representation over the symmetric group, bringing us closer to describing the quasi-invariant polynomials in all characteristics and numbers of variables.
[…]
quasi-invariant polynomials; modular representation theory
[…]
81R12; 13A50; 20C08
@@ 1 Introduction
Let $\mathbbm{k}$ be a field, and consider the action of the symmetric group $S_{n}$ on the space $\mathbbm{k}[x_{1},\dots,x_{n}]$ of $\mathbbm{k}$-valued polynomials by permuting the variables. A polynomial in $\mathbbm{k}[x_{1},\dots,x_{n}]$ is symmetric if it is invariant under this action. Equivalently, since $S_{n}$ is generated by transpositions, a polynomial $K$ is symmetric if $s_{i_{1}i_{2}}K=K$ or $(1-s_{i_{1}i_{2}})K=0$ for all $s_{i_{1}i_{2}}\in S_{n}$. One may consider generalizations of symmetric polynomials in which this condition is relaxed, so that we only require $(1-s_{i_{1}i_{2}})K$ be divisible by some large polynomial. This leads to the notion of quasi-invariant polynomials.
@@
Let $\mathbbm{k}$ be a field. For $m\in\mathbb{Z}_{\geq 0}$, $n\in\mathbb{Z}_{>0}$, a polynomial $K\in\mathbbm{k}[x_{1},\dots,x_{n}]$ is $m$-quasi-invariant if for all $s_{i_{1}i_{2}}\in S_{n}$ we have that $(x_{i_{1}}-x_{i_{2}})^{2m+1}$ divides $(1-s_{i_{1}i_{2}})K$. We denote the space of $m$-quasi-invariants by $Q_{m}(n,\mathbbm{k})$.
@@
Quasi-invariant polynomials were first introduced by Chalykh and Veselov in 1990 [6] to describe the harmonic, zero eigenvalue eigenfunctions of quantum Calogero–Moser systems. Calogero–Moser systems are a collection of one-dimensional dynamical particle systems that were found to be both solvable [4] and integrable [10]. Due to these properties, they have become extensively studied in mathematical physics, with connections to a number of other fields of mathematics, including representation theory.
Quasi-invariant polynomials were also later found to describe the representation theory of the spherical subalgebra of the rational Cherednik algebra [3]. This subalgebra is Morita equivalent to the entire rational Cherednik algebra [7], so quasi-invariants describe representations of rational Cherednik algebras as well. Such algebras have connections to combinatorics, mathematical physics, algebraic geometry, algebraic topology, and more, leading them to become a central topic in representation theory.
@@
In 2003, Felder and Veselov found the Hilbert series of the space of quasi-invariants in characteristic zero [9], proving its freeness in the process. Work on quasi-invariants in characteristic $p$ started in 2020, when Ren and Xu proved a sufficient condition for the Hilbert series of $Q_{m}(n,\mathbf{F}_{p})$ to be different from the Hilbert series of $Q_{m}(n,\mathbf{Q})$ [11]. They accomplished this by computing non-symmetric polynomial “counterexamples” in characteristic $p$, where the polynomial has lower degree than any non-symmetric quasi-invariant polynomial in characteristic 0. They also made several conjectures about quasi-invariants in characteristic $p$, including that the condition they found is also sufficient, the quasi-invariants are free, and that the Hilbert polynomial is palindromic for $p>2$. In 2023, the first author proved a general form for the Hilbert series of the quasi-invariants for $n=3$, $p>3$, proving freeness and the palindromicity of the Hilbert polynomial in the process [13].
We expect the development of the theory of quasi-invariants in characteristic $p$ to be useful in mathematical physics and integrable systems through the theory of $q$-deformed quasi-invariants. These are certain deformations of quasi-invariants in characteristic zero introduced by Chalykh in 2002 [5] used to describe eigenfunctions of Macdonald difference operators, which are a generalization of elliptic Calogero–Moser systems [12]. We expect the theory of quasi-invariants in characteristic $p$ to be related to the theory of $q$-deformed quasi-invariants when $q$ is a root of unity, in analog to the classical connection between representations of Lie algebras in characteristic $p$ and quantized enveloping algebras [2]. We note that a few similarities between these two spaces of quasi-invariants have already been found in [13].
In this paper, we consider the cases $n=3$, $p=2,3$. These cases differ from the $p>3$ case studied in [13] since in $p=2,3$ the representations of $S_{3}$ are modular, i.e., are not completely reducible. Despite these limitations, we describe the Hilbert series explicitly for all $m$, proving the following.
@@
$\mathcal{H}(Q_{m}(3,\mathbbm{k}))=\frac{1+2t^{d}+2t^{6m+3-d}+t^{6m+3}}{(1-t)(1-t^{2})(1-t^{3})},$
where $d=3m+1$ if there is no Ren–Xu counterexample and $d$ is the degree of the minimal degree Ren–Xu counterexample otherwise. In particular, the conditions found in [11] for the Hilbert series of $Q_{m}(3,\mathbbm{k})$ to be different from the Hilbert series of $Q_{m}(3,\mathbf{Q})$ are necessary.
@@
In the case $p=2$, we also define $m$-quasi-invariants in the case where $m$ is a half-integer and prove an analogous statement to Theorem 1.3 in this case. Using quasi-invariants at half-integers, we also compute the generators of $Q_{m}(3,\mathbf{F}_{2})$ as an $\mathbf{F}_{2}[x_{1},x_{2},x_{3}]^{S_{3}}$-module explicitly.
In Section 2, we state some of the basic facts about quasi-invariant polynomials and introduce modular representations of $S_{3}$. In Section 3, we compute the generators of $Q_{m}(3,\mathbf{F}_{2})$, proving Theorem 1.3 for $p=2$ in the process. In Section 4, we begin discussing $p=3$, and show that some properties of quasi-invariants in 3 variables from [13] carry over to the $p=3$ case after converting from the standard representation to the $\mathrm{sign-triv}$ representation. In Section 5, we show that minimal degree Ren–Xu counterexamples are the lowest degree non-symmetric generators for $Q_{m}(3,\mathbf{F}_{3})$ and show that there is one other higher degree generator belonging to the $\mathrm{sign-triv}$ representation. Finally, in Section 6, we consider all other indecomposable representations of $S_{3}$ in $Q_{m}(3,\mathbf{F}_{3})$, finishing the proof of Theorem 1.3 for $p=3$.
@@ 2 Preliminaries
  1. $1$.
    $\mathbbm{k}[x_{1},x_{2},x_{3}]^{S_{3}}\subset Q_{m}(3,\mathbbm{k})$, $Q_{0}(3,\mathbbm{k})=\mathbbm{k}[x_{1},x_{2},x_{3}]$, and $Q_{m}(3,\mathbbm{k})\supset Q_{m^{\prime}}(3,\mathbbm{k})$, where ${m^{\prime}>m}$.
  2. $2$.
@@
Note that [8] proves Proposition 2.1 in the case, where $\mathbbm{k}=\mathbb{C}$. However, the proofs for the first two assertions work over any field, and the last assertion follows from the Hilbert basis theorem. In view of the structure of $Q_{m}(3,\mathbbm{k})$ as a module over the symmetric polynomials, given some $K\in Q_{m}(3,\mathbbm{k})$, we will frequently refer to quasi-invariant polynomials that can be obtained via scalar multiplication of $Q$ by a symmetric polynomial. To distinguish these polynomials from the ordinary $\mathbbm{k}$-multiples of $Q$, we will refer to them as symmetric polynomial multiples of $Q$.
We consider $Q_{m}(3,\mathbf{F}_{2})$ and $Q_{m}(3,\mathbf{F}_{3})$ as representations of $S_{3}$, where $S_{3}$ permutes the variables $x_{1}$, $x_{2}$, $x_{3}$. Since $Q_{m}(3,\mathbf{F}_{2})$ and $Q_{m}(3,\mathbf{F}_{3})$ are vector spaces over $\mathbf{F}_{2}$ and $\mathbf{F}_{3}$ respectively and the characteristics $2$ and $3$ divide $|S_{3}|$, $Q_{m}(3,\mathbf{F}_{2})$ and $Q_{m}(3,\mathbf{F}_{3})$ are modular representations of $S_{3}$.
@@ 2.1 Preliminary definitions for $\boldsymbol{p=2}$
  2. $2$.
    $\mathrm{std}$ is the $2$-dimensional irreducible representation of $S_{3}$ obtained by reducing the standard representation in characteristic 00 mod $2$.
  3. $3$.
    $\mathrm{triv-triv}$ is the $2$-dimensional indecomposable representation that contains a copy of $\mathrm{triv}$ as a subrepresentation such that the quotient of $\mathrm{triv-triv}$ by this subrepresentation is $\mathrm{triv}$.
@@
The polynomial $E_{\mathrm{triv-triv}}:=x_{1}^{2}x_{2}+x_{2}^{2}x_{3}+x_{3}^{2}x_{1}\in\mathbf{F}_{2}[x_{1},x_{2},x_{3}]$ generates a copy of $\mathrm{triv-triv}$. To see this, note that for any $i_{1}$, $i_{2}$, we have
$(1-s_{i_{1}i_{2}})E_{\mathrm{triv-triv}}=x_{1}^{2}x_{2}+x_{1}x_{2}^{2}+x_{1}^{2}x_{3}+x_{1}x_{3}^{2}+x_{2}^{2}x_{3}+x_{2}x_{3}^{2}\in\mathbf{F}_{2}[x_{1},x_{2},x_{3}]^{S_{3}}.$
@@
We cannot define $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv-triv}}$ to exclude copies of $\mathrm{triv}$ since we can add elements of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}}$ to copies of $\mathrm{triv-triv}$ and still obtain a copy of $\mathrm{triv-triv}$. For example, $F:=E_{\mathrm{triv-triv}}+x_{1}^{3}+x_{2}^{3}+x_{3}^{3}$ still satisfies $(1-s_{i_{1}i_{2}})F=(1-s_{i_{1}i_{2}})E_{\mathrm{triv-triv}}$ for all $i_{1}$$i_{2}$, so it generates a copy of $\mathrm{triv-triv}$ by the same argument as Example 2.4.
Proposition 2.6 ([13]).
As an $\mathbf{F}_{2}[x_{1},x_{2},x_{3}]^{S_{3}}$-module, $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}}$ is freely generated by $1$.
Note that by the classification of indecomposables in Proposition 2.3, every extension of $\mathrm{std}$ and every extension of a module by $\mathrm{std}$ splits. Thus $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ is a direct summand of $Q_{m}(3,\mathbf{F}_{2})$ (whose complement is $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv-triv}}$), and we mainly consider $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$. $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ is generated as a $\mathbf{F}_{2}[x_{1},x_{2},x_{3}]^{S_{3}}$-module by homogeneous copies of $\mathrm{std}$, so following [13], we consider generating representations of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ as homogeneous copies of $\mathrm{std}$ in a generators and relations presentation of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ with a minimal generator set.
2.1.1 Quasi-invariants at half-integers
Note that if $\mathbbm{k}$ is a field with $\mathrm{char\,}\mathbbm{k}\neq 2$ and $m\in\mathbb{Z}_{\geq 0}$, then for any $K\in\mathbbm{k}[x_{1},\dots,x_{n}]$, $(x_{i_{1}}-x_{i_{2}})^{2m}|(1-s_{i_{1}i_{2}})K$ implies $(x_{i_{1}}-x_{i_{2}})^{2m+1}|(1-s_{i_{1}i_{2}})K$ since $(1-s_{i_{1}i_{2}})K$ is $s_{i_{1}i_{2}}$-antiinvariant, hence the exponent $12m+$ 1 in the definition of quasi-invariant polynomials. But this does not hold in characteristic 2, since there is no concept of antiinvariants. Indeed, one can check that for $K=x_{1}^{2}+x_{2}^{2}$, we have $(x_{i_{1}}-x_{i_{2}})^{2}|(1-s_{i_{1}i_{2}})K$ for all $i_{1}$, $i_{2}$, but $(x_{i_{1}}-x_{i_{2}})^{3}\nmid|(1-s_{i_{1}i_{2}})K$ if $i_{1}=1,2$$i_{2}\neq 1,2$.
2.2 Preliminary definitions for $\boldsymbol{p=3}$
Proposition 2.7 ([1]).
  1. $1$.
  2. $2$.
  3. $3$.
  4. $4$.
  5. $5$.
  6. $6$.
Provided are some examples of copies of these indecomposable representations:
Example 2.8.
The space $W\subset\mathbf{F}_{3}[x_{1},x_{2},x_{3}]$ spanned by $x_{1}+x_{2}+x_{3}$ and $x_{1}-x_{2}$ over $\mathbf{F}_{3}$ is copy of $\mathrm{sign-triv}$. Indeed, the space $T\subset W$ spanned by $x_{1}+x_{2}+x_{3}$ is a copy of $\mathrm{triv}$. One can check $x_{1}-x_{2}\in W/T$ is acted by negation by all transpositions in $S_{3}$ and $W/T$ is $1$-dimensional so $W/T$ is a copy of $\mathrm{sign}$. Finally, it is easy to show that there are no copies of $\mathrm{triv}$ or $\mathrm{sign}$ in $W$ other than $T$. Since $V$ has a unique irreducible subrepresentation, it is indecomposable, and we conclude that it is a copy of $\mathrm{sign-triv}$.
Example 2.9.
The space $V\subset\mathbf{F}_{3}[x_{1},x_{2},x_{3}]$ consisting of homogeneous linear polynomials is a copy of $\mathrm{triv-sign-triv}$. Indeed, $W\subset V$ from Example 2.8 is a copy of $\mathrm{sign-triv}$. Then $V/W$ is one-dimensional, and one can check that it is a copy of $\mathrm{triv}$. Finally, it is easy to show that there are no copies of $\mathrm{triv}$ or $\mathrm{sign}$ in $V$ other than $T$, so $V$ has a unique irreducible subrepresentation, it is indecomposable, and we conclude that it is a copy of $\mathrm{triv-sign-triv}$.
Example 2.10.
Similarly, one may check that the space $U$ spanned by
$(x_{1}-x_{2})(x_{1}-x_{3})(x_{2}-x_{3}),\qquad-x_{1}^{2}x_{2}-x_{1}^{2}x_{3}+x_{1}x_{2}^{2}+x_{1}x_{3}^{2}$
over $\mathbf{F}_{3}$ is a copy of $\mathrm{triv-sign}$ and that the space spanned by
$(x_{1}-x_{2})(x_{1}-x_{3})(x_{2}-x_{3}),\qquad-x_{1}^{2}x_{2}-x_{1}^{2}x_{3}+x_{1}x_{2}^{2}+x_{1}x_{3}^{2},\qquad(x_{1}-x_{2})x_{1}x_{2}$
is a copy of $\mathrm{sign-triv-sign}$.
Similarly to the $p=2$ case, we define $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ and $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{triv}}$ to be the direct sum of all copies of $\mathrm{sign}$ and $\mathrm{triv}$ in $Q_{m}(3,\mathbf{F}_{3})$, respectively.
Proposition 2.11 ([13]).
As $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$-modules,
  1. $1$.
    $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{triv}}$ is freely generated by $1$.
  2. $2$.
    $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ is freely generated by $\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m+1}$.
Next we define $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ as the direct sum of all copies of $\mathrm{sign}$, $\mathrm{triv}$, and $\mathrm{sign-triv}$. For this paper we consider generators of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ to be homogeneous polynomials other than $1$ and $\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m+1}$ such that they are in the $(-1)$-eigenspace of $s_{12}$ and are in a generators and relations presentation of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ as an $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$-module with the least number of generators. Moreover, if $K$ is a generator of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ then it necessarily generates a copy of $\mathrm{sign-triv}$ since we assumed $K$ neither generates $\mathrm{triv}$ nor $\mathrm{sign}$.
Remark 2.12.
Similar to in the $p=2$ case, we cannot define $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ to exclude copies of $\mathrm{sign}$ since we can add elements of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ to copies of $\mathrm{sign-triv}$ and still obtain a copy of $\mathrm{sign-triv}$. For example, the spaces spanned by
$\big{(}x_{1}^{6}-x_{2}^{6}\big{)}(x_{1}+x_{2}+x_{3})^{3},\qquad\big{(}x_{1}^{6}+x_{2}^{6}+x_{3}^{6}\big{)}(x_{1}+x_{2}+x_{3})^{3}$
and
$\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{3}+\big{(}x_{1}^{6}-x_{2}^{6}\big{)}(x_{1}+x_{2}+x_{3})^{3},\qquad\big{(}x_{1}^{6}+x_{2}^{6}+x_{3}^{6}\big{)}(x_{1}+x_{2}+x_{3})^{3}$
generate two copies of $\mathrm{sign-triv}$ in $Q_{1}(3,\mathbf{F}_{3})$, and their sum contains
$\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{3}\in Q_{1}(3,\mathbf{F}_{3})_{\mathrm{sign}}.$
Remark 2.13.
One could define subspaces of $Q_{m}(3,\mathbf{F}_{3})$ for $\mathrm{triv-sign-triv}$, $\mathrm{sign-triv-sign}$, $\mathrm{triv-sign}$ similar to $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$, however this is not particularly helpful, as unlike for ${p=2}$, we cannot decompose $Q_{m}(3,\mathbf{F}_{3})$ into a direct sum of subspaces of this form. The space $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is still relevant, as it is the critical piece to understanding quasi-invariants in characteristic 3, as we see in Sections 4 and 5.
3 Quasi-invariants in characteristic 2
In this section, we write down explicit generators for $Q_{m}(3,\mathbf{F}_{2})$ and prove Theorem 1.3 for $p=2$. Note that we already know the structure of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}}$ from Proposition 2.6. We start by extending this to $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv-triv}}$.